calculate the slope of the line: x-4x=3
Found 2 solutions by nerdybill, Fombitz:
Answer by nerdybill(7384)

Put
x-4x=3
into the "slope-intercept" form
y=mx+b
where
m is the slope
b is the y-intercept
.
x-4x=3
-4x = -x+3
x = (1/4)x - 3/4
Slope then is 1/4

Answer by Fombitz(32388)

Is your equation correct?
If so, it's only a function of x and would have infinite slope.
In other words, it would be a vertical line.
If that's not what you meant, then please re-post.